Six pairs of data yield r = 0.444 and the regression equation y hat = 5x + 2. Also, y bar = 18.3. What is the best predicted value of y for x = 5?
Answer #1
Concepts and reason

The concept regression and predicting the value using the regression equation are used here.

A regression equation is a linear equation which is used to predict a dependent variable based on the independent variable. This regression equation represents the relationship between the dependent and the independent variable.

Fundamentals

The regression equation between dependent variable, yy and the independent variable, xx can be represented by the following equation:

y^=a+bx\hat y = a + bx

In the above regression equation, the values aa and bb are the intercept and slope coefficient of the equation respectively.

The above regression equation can be used to predict the yy for future xx as follows:

y^(pred)=a+bx(future)\hat y\left( {{\rm{pred}}} \right) = a + bx\left( {{\rm{future}}} \right)

The “pred” represents the predicted and future represents the new values of xx . But the prediction of the future value will be risky outside the pattern of the observed value.

The best predicted value of yy for future xx is,

y^(pred)=yˉ\hat y\left( {{\rm{pred}}} \right) = \bar y

The value of r is provided as 0.444 which indicates that there is a weak relationship between the variable x and y.

The regression equation is provided as y^=5x+2\hat y = 5x + 2 and the value of yˉ\bar y is 18.3.

The best predicted value of y for x=5x = 5 can be computed as:

y^(pred)=yˉ=18.3\begin{array}{c}\\\hat y\left( {{\rm{pred}}} \right) = \bar y\\\\ = 18.3\\\end{array}

Ans:

The best predicted value of y for x=5x = 5 is 18.3.